OpenStudy (anonymous):

Fractional Equation help please? I will put the equation in the comments below.

OpenStudy (anonymous):

\[\frac{ 1 }{ y^2-y-2 }-\frac{ 2 }{ y-2 }=\frac{ 1 }{ y+1 }\]

OpenStudy (anonymous):

I got y=-1 but that wasn't one of the choices...

OpenStudy (jdoe0001):

so.. if you factor the 1st denominator, the trinomial that is what factors does that give you anyhow?

OpenStudy (anonymous):

(y-2)(y+1)

OpenStudy (jdoe0001):

hmm lemme ... right

OpenStudy (anonymous):

So I multiply each fraction by (y-2)(y+1)?

OpenStudy (anonymous):

Leaving me with \[1-2(y+1)=y-2\]

OpenStudy (anonymous):

Then I distribute the -2 into y+1 and get 1-2y-2=y-2

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

Then I simplify to get 1=3y so y is 1/3 and wow i just solved it -_- ok i know what i was doing wrong XD

OpenStudy (anonymous):

lol thanks anyway

OpenStudy (jdoe0001):

hehe